### v2

# Benchmarking Deep Learning Methods for Irradiance Estimation from Sky Images with Applications to Video Prediction-Based Irradiance Nowcasting

Abstract
To address the high levels of uncertainty associated with photovoltaic energy, an increasing number of studies focusing on short-term solar forecasting (i.e. nowcasting) have been published. Most of these studies use deep-learning-based models to directly forecast a solar irradiance or photovoltaic power value given an input sequence of sky images. Recently, however, advances in generative modeling have led to approaches that divide the nowcasting problem into two sub-problems: 1) future event prediction, i.e. generating future sky images; and 2) solar irradiance or photovoltaic power estimation, i.e. predicting the concurrent value from a single image. One such approach is the SkyGPT model, whose potential for improvement is shown to be much larger in the estimation component than in the generative component. Thus, in this paper, we focus on the solar irradiance estimation problem and conduct an extensive benchmark of deep learning architectures across the widely-used Folsom, SIRTA and NREL datasets. Moreover, we perform ablation experiments on different training configurations and data processing techniques, including the choice of the target variable used for training and adjustments of the timestamp alignment between images and irradiance measurements. In particular, we draw attention to a potential error associated with the sky image timestamps in the Folsom dataset and suggest a possible fix. By leveraging the three datasets, we demonstrate that our findings are consistent across different solar stations. Finally, we combine our best irradiance estimation model with a video prediction model and obtain state-of-the-art results on the SIRTA dataset.
keywords:
Solar Irradiance , Nowcasting , Forecasting , Deep Learning , Sky Images , Video Prediction
journal: Solar Energy
[…]
[label1, label2]organization=Department of Electrical and Electronic Engineering, Federal University of Santa Catarina,addressline=, city=Florianópolis, postcode=, state=Santa Catarina, country=Brazil
1 Introduction
Solar photovoltaic (PV) energy, which harvests the Sun’s energy directly into electric energy by using semiconductor cells, has seen an exponential growth in recent years [1]. This has brought forth a new set of challenges to the electrical grid operators [2, 3, 4, 5], due to the intermittent and uncertain nature of this energy source [6, 7]. The primary cause of uncertainty in PV power generation is the presence of clouds, which can abruptly increase or decrease the amount of solar radiation harnessed by PV panels at any given time. Consequently, a large number of studies have been published in recent years with the goal of creating accurate solar forecasting models.
In the very short term solar forecasting ($\leq$ 60 minutes forecasting horizon), commonly referred to as solar nowcasting, all-sky image (ASI) based models have achieved state-of-the-art results, due to their ability to capture the cloud information with a very high spatio-temporal resolution [8]. These images are typically combined with convolutional neural networks (CNN), which extract features from the images that will be used to predict the solar value. The following paragraphs briefly describe some of these approaches.
The authors of [9] use a ResNet18 architecture [10] to encode the image features, which are then used to predict the solar irradiance 5 to 10 minutes ahead. However, instead of feeding a static image to the model, they stack multiple images at different timestamps and gain up to a 7% performance boost. The method proposed in [11] used a 3D CNN [12] to forecast the clear sky index $k_{t}$, which results in a 15.2% improvement over the smart persistence model. In [13], a benchmark study of four different architectures (CNN, CNN+LSTM, 3D-CNN and ConvLSTM) is conducted and, while the more complex models tend to perform better, it is shown that they all share a fundamental problem: the predicted future irradiance values are often delayed relative to the ground truth measurements, resulting in irradiance ramp predictions that are systematically delayed relative to the ground truth.
A multi-modal transformer based framework is proposed in [14], where the solar irradiance time series is encoded by a transformer [15], and the sky images are transformed into optical flow maps and encoded by a vision transformer [16], which are then fused with a cross modality attention block to predict the solar irradiance up to 30 minutes ahead. In [17], a VGG16 [18] inspired model is developed specifically for the solar irradiance nowcasting task, the SolarNet model, which outperformed several of their benchmark models. In [19], different input and output configurations are explored for the Sunset model, which is another model specifically designed for the solar nowcasting task, but focuses on the solar PV power instead of solar irradiance.
The methodology proposed in [20] is particularly interesting as it separates the future event prediction task from the irradiance estimation task. This is achieved by using a sky image generation model, SkyGPT, which is trained auto-regressively to predict future sky images based on past image sequences. With the generated future sky image, a modified version of the U-Net architecture [21] is then used to estimate the concurrent PV power. By decoupling the future event prediction task (SkyGPT) from the PV power estimation task (U-Net), they demonstrate that improving the SkyGPT model (such that it could perfectly predict the future sky image) would only boost the performance by 13%, whereas improving the estimation model could yield up to a 64% performance boost (see Figure 7 of [20]).
Inspired by the larger potential improvement, this study focuses primarily on the solar irradiance estimation task, using exclusively sky images as input, i.e. given an image at timestamp $t$, we want to predict the corresponding irradiance value at the same timestamp $t$. Although this is different from PV power estimation, we expect the results to generalize well across both tasks, since they share the same underlying nature and are highly correlated.
Compared to solar nowcasting, fewer studies have tackled the solar estimation task. In [22] they perform solar irradiance estimation and show that appending a Sun mask as a fourth channel to the images significantly increases the estimation performance, which we also investigate in this study. In [23] they adapt the Sunset model for the solar PV power estimation task, which we use in this study for irradiance estimation.
We begin our study by conducting an extensive irradiance estimation benchmark of ten deep learning architectures across the Folsom, SIRTA and NREL datasets, which are three datasets commonly used in the solar forecasting literature. We then perform several ablation experiments, including the choice of the target variable used for training and adjusting the timestamp alignment between the images and irradiance measurements. Finally, we train a video prediction model and combine it with our irradiance estimation model to conduct proof of concept experiments on the two-step approach in solar irradiance nowcasting. All of our experiments are evaluated using the root mean squared error (RMSE) and mean absolute error (MAE) and we show that our findings are consistent across all three datasets. Our main contributions can be summarized as follows:
  1. 1.
    We conduct an extensive irradiance estimation benchmark of deep learning architectures across three solar irradiance datasets and find that the ResNet50 architecture provides significant gains over the U-Net architecture used in [20];
  2. 2.
    We demonstrate that appending a Sun mask as a fourth channel to the images as in [22] improves the performance when the model’s target variable is the unnormalized solar irradiance value, but is detrimental otherwise;
  3. 3.
    We demonstrate that the Folsom dataset contains an error associated with their sky image timestamps and a potential fix is investigated;
  4. 4.
    We demonstrate that shifting the irradiance measurements of the training set by a few seconds in time can significantly improve the test set performance.
  5. 5.
    With our best irradiance estimation model and a simple video prediction model, we obtain state of art nowcasting results on the SIRTA dataset.
The rest of this article is structured as follows: section 2 briefly describes the fundamental concepts of solar irradiance, such as the different irradiance components and clear sky models; section 3 describes the three datasets that were used in this study, as well as the different processing steps that were applied; section 4 describes our methods, where we go into detail about our trained models and evaluation metrics; section 5 details all the results and experiments conducted in this study; section 6 concludes the article.
2 Solar Irradiance Fundamentals
2.1 Irradiance Components
Solar irradiance is a measurement of the rate of solar radiation energy that arrives at a surface area and is typically measured in $\text{W/m}^{2}$. The mean solar irradiance at the top of the Earth’s atmosphere, also known as the solar constant [24], is approximately 1366 $\text{W/m}^{2}$. After entering the Earth’s atmosphere, the solar radiation will be attenuated by the complex interactions with the atmospheric constituents, resulting in two irradiance components at the ground level: diffuse horizontal irradiance (DHI) and direct normal irradiance (DNI). The DHI is the amount of solar radiation incident upon a horizontal plane that has been scattered or reflected by atmospheric constituents or nearby objects. The DNI measures the amount of solar radiation that is coming directly from the direction of the Sun beam. The sum of these two components on a horizontal plane results in the global horizontal irradiance (GHI), denoted by:
$\text{GHI}=\text{DNI}\times\cos{\theta_{z}}+\text{DHI}$ (1)
where $\theta_{z}$ is the Solar Zenith Angle (SZA). This component is the primary focus of this study.
2.2 Clear Sky Models
Predicting the GHI under clear sky conditions is a well-defined problem, and there are hundreds of different clear sky models available in the literature, with varying but generally excellent performances across many different climates [25, 26]. Clear sky models can also be used to obtain the clear sky index $k_{t}$, that normalizes the GHI measurement according to:
$k_{t}=\frac{I}{I^{\text{clr}}}$ (2)
where $I$ is the GHI measurement and $I^{\text{clr}}$ is the clear sky GHI predicted by a clear sky model. In this study, the Simplified Solis Clear Sky Model [27] is used.
Intuitively, $I^{\text{clr}}$ would represent the highest possible GHI value and $k_{t}$ should range between 0 (overcast) to 1 (clear sky). However, $k_{t}$ can assume values greater than 1, not only because of the imperfections of the clear sky model, but mainly due to a phenomenon called overirradiance [28], where the presence of clouds actually increases the GHI through solar beam reflection.
Other than normalization, the $k_{t}$ index and the clear sky prediction $I^{\text{clr}}$ can be used to obtain the smart persistence model (SPM), which is a naive model that assumes that the current $k_{t}$ will stay the same throughout the entire forecasting horizon $h$:
$I_{t+h}=k_{t}\times I^{\text{clr}}_{t+h}$ (3)
This model is commonly used as a baseline reference for solar forecasting models.
Another commonly used index is the clearness index $K_{t}$ [29], which uses the extraterrestrial irradiance instead to normalize the GHI:
$K_{t}=\frac{I}{I^{\text{extr}}}$ (4)
where $I^{\text{extr}}$ is the extraterrestrial GHI. The main advantage of this index is that $I^{\text{extr}}$ is much easier to model than $I^{\text{clr}}$, and can be written as:
$I^{\text{extr}}=I_{0}\times\cos{\theta_{z}}$ (5)
where $I_{0}$ is the solar constant and $\theta_{z}$ is the SZA. The choice between both normalization indices is investigated in subsubsection 5.2.3.
2.3 Detection of Clear Sky Periods
Although sky condition classification is outside the scope of this study, detecting clear/cloudy sky periods can still be used to split the testing set into two smaller subsets, thus providing a more detailed and insightful evaluation of the model’s performance under varying sky conditions. Based on a recent review of 21 different clear sky identification methods [30], and considering the fact that the algorithm is readily available through Python’s pvlib library, we chose the Reno [31] method. This is a very simple and effective algorithm that only requires the measured GHI and the expected clear sky GHI data (from the Simplified Solis model, in our case) as inputs. By comparing 5 different statistics of the measured and expected clear sky values over a sliding time window (typically 10 minutes), the algorithm flags a measurement as either clear or cloudy.
Figure 1 shows an example of this method being applied. From our testing, this method works very well for all three datasets in the majority of cases. For more details on this algorithm, we refer the reader to the original paper [31].
Figure 1: Clear and cloudy sky measurements in each dataset obtained through the Reno algorithm [31].
3 Data Description and Processing
In this section, we describe the raw data used in this study, obtained from three different datasets, as well as the preprocessing pipeline that is made prior to training the models.
3.1 Folsom Dataset
The first dataset used in this study is the Folsom dataset ($\mathcal{D}^{\text{folsom}}$) [32], collected in Folsom, California (38.642 N and 121.148 W), over a 3-year period (2014-2016). This dataset has been widely used by the solar forecasting community [22, 9, 33, 34, 35, 36, 37, 38] and it contains a variety of irradiance correlated data, such as sky images, temperature, wind, humidity, among others. In this benchmark study, only the sky images and irradiance data were used.
The sky images were captured during the daylight by a fish-eye lens camera with a resolution of $1536\times 1536$ pixels at a 1-minute sampling interval, totaling over 750,000 images. The GHI and DHI data are collected throughout the entire day by a second generation RSR, which contains two Licor LI-200SZ pyranometers, and the DNI is computed from the measured GHI, DHI and the solar zenith angle $\theta_{z}$, as per Equation 1. While all three irradiance data are measured at a 1-second sampling interval, only their 1-minute averages are provided.
Examples of the sky images are shown in the leftmost column of Figure 2 and, unlike the other two datasets, the camera exposure time is the same for all images in $\mathcal{D}^{\text{folsom}}$, though the exact value of this setting is unavailable in [32] and the image metadata. Also unlike the other datasets, $\mathcal{D}^{\text{folsom}}$ exhibits a few instances of lens contamination due to dust, dirt, and condensation. Although we believe that automatic mitigation strategies for this issue are beyond the scope of this study, we think it is important to highlight it and to inform future research, since this can have an impact on model performance. Thus, we manually identified three representative examples in Figure 3. The GHI distribution of $\mathcal{D}^{\text{folsom}}$ is shown in the left panel of Figure 4, where it is clear that the density of large irradiance values in this dataset is much higher, i.e. it has a larger number of clear sky instances than the other two datasets.
3.2 SIRTA Dataset
The second dataset we explored was the SIRTA dataset ($\mathcal{D}^{\text{sirta}}$) [39], which was collected in Palaiseau, France (48.713 N and 2.208 E) at the Site Instrumental de Recherche par Télédétection Atmosphérique (SIRTA) observatory. The SIRTA observatory has been rigorously collecting high quality irradiance, sky image and meteorological data for more than 10 years and their dataset is also extensively used by the community [13, 40, 41, 42, 43, 44, 45]. For our benchmark study, we only select the 2017-2019 data, as this seems to be the most common choice across the literature.
The sky images (middle images of Figure 2) were captured with a $768\times 1024$ resolution at a 1-minute sampling interval in 2017 and 2-minute in 2018 and 2019 and come in two forms: 1) a long exposure form that shows the sky in more detail but can have saturated pixels in the area around the Sun (1/1000 s exposure); 2) a short exposure form that shows the area around the Sun less saturated and in more detail, but the rest of the sky is shown in less detail (1/2000 s exposure). The GHI and DHI are measured by a CMP22 pyranometer and the DNI is measured by a CHP1 pyrheliometer, all at a 1-second sampling interval, but only the 1-minute averages are provided.
Compared to the other two datasets, the GHI distribution in $\mathcal{D}^{\text{sirta}}$ (middle panel of Figure 4) is much more concentrated at lower values, i.e. this dataset has much more cloudy/overcast samples which, generally speaking, are much harder for the models to learn than clear sky samples.
3.3 NREL Dataset
Lastly, we also analyzed the National Renewable Energy Laboratory (NREL) dataset ($\mathcal{D}^{\text{nrel}}$) [46], collected at NREL’s Solar Radiation Research Laboratory (SRRL), located in Golden, Colorado (39.742 N and 105.18 W). The SSRL has been measuring irradiance correlated data since 1981 and, once again, many studies have been published using this dataset: [47, 48, 14, 49, 50]. For our benchmark studies, we only selected the sky images and irradiance data collected from 2020 to 2023. Unlike the other datasets, $\mathcal{D}^{\text{nrel}}$ provides multiple independent measurements of the same irradiance component. We used the following sensors:
  1. 1.
    DNI: CHP1-1, CHP1-2, sNIP, NIP and DR02 pyrheliometers;
  2. 2.
    DHI: CM22-2, CM22-1, SR25, Research 1 and 8-48 pyranometers;
  3. 3.
    GHI: CMP22, CMP22-1, CMP22-2, CMP11 and CM6b pyranometers.
By leveraging all of these sensors, we were able to detect and filter out several faulty measurements. To do so, we implement the following: let $M(X)$ denote the median of a set of measurements $X$. We first compute the medians of the DHI and DNI measurements, $M(\text{DHI})$ and $M(\text{DNI})$, and use them to estimate the global component, $\text{GHI}_{\text{calculated}}$, according to Equation 1. We then compare this calculated value with the median of the GHI measurements, $M(\text{GHI})$, and keep only the samples that satisfy:
$|\text{GHI}_{\text{calculated}}-M(\text{GHI})|\leq 10\text{ W/m}^{2}$ (6)
This filtering step retains 961,069 of the original 1,063,282 measurements and improves the overall reliability of the irradiance labels. The GHI distribution is plotted in the right panel of Figure 4.
The sky images are provided at a $1536\times 1536$ pixel resolution in both long and short exposure times (rightmost images of Figure 2), though the exact value of this setting is also unavailable in this dataset. The sky images are only available once every 10 minutes, which makes this a much smaller dataset compared to the other two (for our selected years in each dataset). The impact that a smaller training dataset has on the model performance is further discussed in subsubsection 5.2.5
Figure 2: Clear and cloudy sky image examples for each dataset used in this study. $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ also provide a short exposure alternative for each image, so that the area around the Sun can be shown in greater detail. The depicted images were cropped such that they have the same shape.
Figure 3: Lens contamination examples in $\mathcal{D}^{\text{folsom}}$: an image with severe contamination (left), an image taken immediately before lens cleaning (middle), and one taken immediately after cleaning (right).
3.4 Sky Image Data Processing
The image processing was essentially the same for all 3 datasets. First we compute the mean image of each dataset (using only the long exposure images for $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$) and then define a circular region of interest (ROI), with radius $R$ and center $C$, which is used as a binary mask for that dataset (see Table 1). These masks are applied to all the images in their respective datasets to erase every pixel outside their ROI. This is done with the intention of standardizing the image data across all datasets, such as removing the timestamp annotation in the upper left corner of the $\mathcal{D}^{\text{nrel}}$ images and removing any lens flare and/or noise effect in the black area of the images. Following this, we center square crop the images on $C$ with length $2R$. Finally, the images are downscaled to a $64\times 64$ resolution using the Python’s PIL library. To avoid any potential JPEG compression losses [51], the resized images are saved to disk in .npy format.
Figure 4: The GHI data distributions for the three datasets used in this study, plotted after the data processing pipeline. This distribution is based on the total number of image/irradiance pairs, not the total number of irradiance measurements.
Table 1: Circular region of interest (ROI) radius and center values used in each dataset.
Dataset $R$ (pixels) $C$ (x, y)
Folsom 755 (768, 768)
SIRTA 340 (502, 394)
NREL 755 (768, 768)
3.5 Irradiance Data Processing
The irradiance data processing was also largely the same for all 3 datasets. First, we resample the data to a 1-second frequency through linear interpolation (more details on why this is done in the following section). To avoid large interpolation errors, this is only done for the samples that are 1-minute apart, i.e. periods with 2 or more minutes of missing data are discarded. Following the linear interpolation, we remove all samples where $\theta_{z}>80^{\circ}$, as these samples provide little information for our application.
The next processing step is only applied to the training set of each dataset (see subsection 3.8 for the train/test splits) and consists of shifting the irradiance measurements in time by applying a time offset ($\Delta\text{t}$), which we do according to:
$T_{\text{new}}=T_{\text{orig.}}+\Delta\text{t}$ (7)
where $T_{\text{new}}$ and $T_{\text{orig.}}$ are the new and original irradiance measurement timestamps, respectively. If a $\Delta\text{t}<0$ is applied, then we are advancing the irradiance series in time, relative to the sky image series, whereas a $\Delta\text{t}>0$ would have the opposite effect, as shown in Figure 5. Depending on how these irradiance values were averaged, we hypothesize that this can be beneficial. For instance, if the value of $I$ at 14:00:00 represents the average GHI values from 13:59:01 to 14:00:00 (backward average), then a $\Delta\text{t}<0$ could diminish the delay introduced by such an average. Thus, we shift the irradiance measurements by -20s, 10s and 30s in $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively, but also investigate other time offsets in subsubsection 5.2.2. The reason why a $\Delta\text{t}$ is not applied to the testing set is because doing so would lead to unfair comparisons among the nowcasting results. For example, if $\Delta\text{t}>0$ (e.g. 30s) is applied to the testing set, then, given the same ASI context, the evaluation of a 10 minute lead time prediction in the shifted time series would be equivalent to a 9:30 minutes lead time in the unshifted time series, which is fundamentally easier to predict. Similarly using $\Delta\text{t}<0$ (e.g. -30s) would be equivalent to a 10:30 minute lead time, which is fundamentally harder.
Figure 5: The effect of applying a $\Delta\text{t}$ to the irradiance time series. A negative $\Delta\text{t}$ essentially advances the samples in time, by shifting the curve to the left, whereas a positive $\Delta\text{t}$ has the opposite effect, by shifting it to the right.
3.6 Sky Image and Irradiance Alignment
Our sky image and irradiance alignment for each dataset was very different, especially for $\mathcal{D}^{\text{folsom}}$. Initially, we aligned the sky images with the irradiance data based on the name of the image files, i.e. an image sample $x^{\text{folsom}}_{t}$, with file name timestamp $t$, was labeled with the irradiance sample $I^{\text{folsom}}_{t}$. While this seems like the most natural choice, we noticed, during our initial experiments, that the model’s largest errors were samples were the prediction actually seemed quite reasonable, since it predicted larger GHI values when the Sun was clearly visible and lower values when it was not. This raised the suspicion that the labels might be misaligned. To the best of our knowledge, this unexpected behavior in $\mathcal{D}^{\text{folsom}}$ has not yet been discussed in the literature and, upon further investigation, we found that the samples taken during time windows with fast moving cloud conditions (e.g. Figure 6) corresponded to the highest errors, leading us to discover a discrepancy between the date modified and the image file name timestamps. In Figure 6 we show two distinct 5 minute GHI/DNI windows in the middle row, with the image date modified alignment in the top row and the image file name alignment in the bottom row. Both cases show that the GHI began at a medium-to-low value, which indicates some cloud coverage, before rapidly increasing, which indicates loss of cloud coverage and a more visible Sun. This behavior seems much more consistent with the top row images, strongly suggesting that it is actually the date modified timestamps that correctly correspond to the GHI time series. However, due to the strong presence of clouds, one could argue that this increase in GHI could be due to an extreme overirradiance event. However, since the DNI component is also being plotted, we can rule out this possibility and confirm that the Sun should indeed transition from covered to visible during both time windows.
Figure 6: Two distinct 5-minute GHI and DNI windows in $\mathcal{D}^{\text{folsom}}$ which show the corresponding Date Modified (DM) aligned images in the top row and the File Name (FN) aligned images in the bottom row. The middle row shows the linearly interpolated GHI data, plotted in UTC-8 timezone. The titles above each image indicate their corresponding FN and DM timestamps (converted to UTC-8 timezone), as well as their JPEG file names. The * marker indicates the timestamp used to align the image to the irradiance plot.
In the 5-minute windows being shown in Figure 6, the file name is approximately 11 minutes 30 seconds ahead of the date modified, but, from our analysis, we found that this difference changes throughout the dataset, as depicted in Figure 7, which clearly indicates a time drift between both timestamps [^1] We hypothesize that the file name timestamps are based on the camera’s internal clock, while the date modified is based on the database’s clock. Assuming the irradiance timestamps are synchronized with the latter, then the file name timestamp would indeed result in misaligned image-irradiance pairs.. Since the date modified seems much better aligned with the GHI time series, we use it to label the images in $\mathcal{D}^{\text{folsom}}$, except during the experiment in subsubsection 5.2.1, where we provide a more quantitative analysis on this issue.
Figure 7: Daily average difference between the file name and date modified of the image data in $\mathcal{D}^{\text{folsom}}$. A positive value indicates that the file name is ahead of the date modified and vice-versa.
As a precautionary measure, we also compared the metadata of $x^{\text{sirta}}$ and $x^{\text{nrel}}$ to their file names, and found that the NREL images also contain an irregularity. However, the date modified issue in this dataset seems much more complex than in $\mathcal{D}^{\text{folsom}}$. In some cases, aligning the images with the date modified yielded a more sensible irradiance label, while in others, the opposite was true. This led us to choose the most simple solution possible, in which we discard all samples where $|\text{Timestamp}(\text{Date Modified})-\text{Timestamp}(\text{File Name})|>30\text{ s}$. This processing step retains 65,202 of the original 66,908 image/irradiance pairs and ensures a more proper alignment for all retained pairs.
Another thing to keep in mind when aligning the images with the irradiance data is the irregularity in the image timestamps, which is especially true assuming the date modified as the correct timestamp (see Figure 8). Since the irradiance samples are all synchronized at :00 seconds, aligning an image that was captured at :30 seconds with the nearest irradiance sample could result in a large error, especially in fast-moving cloud conditions. To mitigate this problem, we linearly interpolate the irradiance samples before the alignment. While this mostly affects $\mathcal{D}^{\text{folsom}}$, the linear interpolation is applied to the other two datasets as well.
Figure 8: Image timestamp second counts in $\mathcal{D}^{\text{folsom}}$. The upper axis shows that the file name timestamps are mostly synchronized at either :$0$ or :59 seconds (i.e. the file ends in “$0$.jpg” or “59.jpg”). On the other hand, the lower axis shows that the date modified timestamps are uniformly spaced throughout the 60 possible values.
3.7 Sun Mask Generation
In subsubsection 5.2.6 we conduct an ablation experiment where we use a Sun mask as an additional input to the estimation model, which has shown to significantly benefit the performance in [22]. First, we compute the distance of the Sun center from the center of the image plane, denoted by R, which is a function of the focal length of the camera f and the angle from the optical axis $\Phi$. This mapping function will vary, since different types of lenses will have different projection types [52]. For the stereographic projection, the mapping function is $\text{R}=2f\tan{\frac{\Phi}{2}}$, and for the equidistant projection the mapping function is $\text{R}=f\Phi$, where $\Phi$ is replaced by the zenith angle $\theta_{z}$. Following this, we compute the Sun center coordinates on the image plane ($\text{x}_{c}$, $\text{y}_{c}$) according to:
$\text{x}_{c}=\text{R}\sin(\theta_{a}-\theta_{c})$ (8)
$\text{y}_{c}=\text{R}\cos(\theta_{a}-\theta_{c})$ (9)
where $\theta_{a}$ is the azimuth angle and $\theta_{c}$ is a correction angle that compensates for the orientation of the camera. To transform $\text{x}_{c}$ and $\text{y}_{c}$ to pixel coordinates and obtain $\text{x}_{p}$ and $\text{y}_{p}$, we apply:
$\text{x}_{p}=\frac{W}{2}(1+\text{x}_{c})$ (10)
$\text{y}_{p}=\frac{W}{2}(1+\text{y}_{c})$ (11)
where $W$ is the image width. Finally, we create the binary Sun mask as a circle with a 5 pixel radius (for $W=64$) and centered on $\text{x}_{p}$ and $\text{y}_{p}$, as shown in Figure 9.
In the original paper, they only apply this mask in $\mathcal{D}^{\text{folsom}}$, in which they determine the stereographic projection for the R mapping function, $f=0.48$ and $\theta_{c}=165^{\circ}$. However, using these same parameters in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ yielded bad results, so, through trial and error, we tuned them according to Table 2. The other two parameters, $\theta_{z}$ and $\theta_{a}$, are obtained with Python’s pvlib library. For more details on this algorithm, we refer the reader to the original paper [22].
Figure 9: Sun mask examples applied to the 64 $\times$ 64 images in all three datasets under cloudy and clear sky conditions. For a 64 $\times$ 64 image, the circle has a 5 pixel radius and is centered on $\text{x}_{p}$ and $\text{y}_{p}$.
Table 2: Sun center localization parameters for each dataset. f is the focal length and $\theta_{c}$ is the correction angle.
Dataset Projection f $\theta_{c}$
Folsom Stereographic 0.48 165
SIRTA Equidistant 0.63 182
NREL Equidistant 0.67 180
3.8 Train, Test and Validation Splits
All three datasets are first split into a train set, which is used to optimize the parameters of the models, and a testing set, which is used to evaluate said models. To account for all possible seasonal variations, we select one full year for testing, which was 2016, 2019 and 2023 for $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively. Moreover, the algorithm described in subsection 2.3 is used to further split the testing set into two additional sets: a clear sky and a cloudy sky set.
The remaining years of each data set are used for the train set. In order to optimize the hyperparameters of the model, we split the full train set into a smaller train set and a validation set using $K$-Fold cross validation (CV). More specifically, we use stratified group $K$-Fold (with $K=5$), where each day is assigned to a different group (to avoid leaking samples from the same day in the training/validation sets) and each bin in the original train set (see Figure 4) is assigned to a different class to maintain the original irradiance distribution in all 5 folds. After optimizing the hyperparameters of each model, we use the full train set to train a final model, which will be evaluated on the test set. Table 3 shows the number of samples and year of each set, after the application of all the pre-processing steps.
Table 3: Number of samples and year of each set after the application of all the pre-processing steps.
Train Test
Dataset All Year All Clear Cloudy Year
Folsom 426,515 2014-2015 220,755 117,423 103,332 2016
SIRTA 297,774 2017-2018 103,926 18,447 85,479 2019
NREL 48,329 2020-2022 16,873 6,973 9,900 2023
4 Methods
Our study investigates a two-step forecasting approach, akin to SkyGPT [20], depicted in Figure 10. A compelling aspect of this approach is that both modules can be designed and optimized entirely independently, i.e., the approach completely decouples future event prediction and GHI estimation. Furthermore, this approach allows us to obtain a lower bound on the GHI forecasting error, simply by replacing the predicted future images with the ground truth future images. Our experiments target primarily the GHI estimation module, however we also include proof of concept experiments where we compare the two-step and single-step forecasting approaches depicted in Figure 10.
In subsection 4.1, we detail our training setup, the model architectures we trained and the evaluation metrics used in the GHI estimation experiments. All of the GHI estimation models are trained and evaluated separately for each dataset, however we also evaluate the aggregated performance across all three datasets.
In subsection 4.2, we briefly describe our training objectives and implementations details for the forecasting experiments, which are only evaluated on $\mathcal{D}^{\text{sirta}}$.
Figure 10: Two-step and single-step approaches for GHI nowcasting. Given an ASI sequence input of $T$ frames, each frame is transformed into a sequence of $N$ non-overlapping patches. In the two-step approach, these patch sequences will be fed into a video prediction module that predicts future ASI frames, and a subsequent module estimates the corresponding irradiance value for each generated frame. It is important to note that this irradiance estimation module is trained independently on the ground truth images. In the single-step approach, the irradiance forecasting module directly predicts the future irradiance values, by applying a linear regression head to the output of the learnable patch embedding, which we prepend to all $T$ patch sequences.
4.1 GHI Estimation
4.1.1 Training Setup
Given an input image $x_{i}$, each model produces an estimate $\hat{I}_{i}$ of the concurrent GHI measurement $I_{i}$. The models are trained to minimize the Mean Squared Error (MSE) loss function given by:
$L_{\text{MSE}}=\frac{1}{B}\sum_{i=1}^{B}(y_{i}-\hat{y}_{i})^{2}$ (12)
where $B$ is the batch size, $y_{i}$ is the ground truth label of the $i^{th}$ sample and $\hat{y}_{i}$ is the model’s output for the $i^{th}$ sample. Following previous work [11, 42], we use the clear sky index as our default target variable, $y_{i}=k_{t,i}$, which implies that the GHI estimate is computed as $\hat{I}_{i}=I^{\text{clr}}\hat{k}_{t,i}$, where $\hat{k}_{t,i}=\hat{y}_{i}$. Other target variables and loss functions are explored in subsubsection 5.2.3.
Our models are trained with a learning rate that exponentially decays from its initial value $\text{lr}_{0}$ to $\frac{\text{lr}_{0}}{10}$ during the first 75% of training according to:
$\text{lr}_{i}=\text{lr}_{i-1}\times e^{\frac{\ln{1/10}}{0.75\times\text{n\_epochs}}}$ (13)
where n_epochs is the total number of epochs and $0<i<0.75\times\text{n\_epochs}$. The remaining 25% of training is done with the learning rate fixed at $\frac{\text{lr}_{0}}{10}$ and we start averaging the weights of the model at the end of every epoch, as this has shown to lead to better generalization [53]. Each model is trained on a NVIDIA GeForce RTX 3090 GPU (24 GB memory) for a total of 16 epochs, except during hyperparemeter optimization, where they are only trained for 8 epochs to save on computational time. Details on the initial learning rate and other hyperparameters values are described in the following section.
4.1.2 Model Architectures and Hyperparameter Optimization
We train and compare 10 different model architectures in this study, 7 of which are popular deep learning architectures used in many computer vision tasks and 3 that were designed/adapted specifically for the solar task. The first 7 include: ResNets (18, 34 and 50) [10]; VGG16 [18]; EfficientNetV2-S [54]; RegNetY-1.6GF [55]; MobileNetV2 [56]. We use the Pytorch implementation of these models and always initialize them with their pre-trained weights, except for the final layer, since we change it to only predict a single output, corresponding to the irradiance value. The other 3 models already output a single value, since they were designed for the solar task, however they are randomly initialized since we were unable to access their pre-trained weights. They are: SolarNet [17]; Sunset [23, 19]; and a modified version of the popular U-Net [21] architecture, modified in [20].
We search 30 different hyperparameter combinations per model per dataset, totaling 900 different searches, which are made through Bayesian optimization [57] using the train and validation splits defined in subsection 3.8 and with the goal of minimizing the average validation loss across all 5 folds. For hyperparameter optimization, we reduce the number of training epochs to 8 and only use 10% of the training data in each dataset. The hyperparameter search space is shown in Table 4.
Table 4: Hyperparameter search space for all the models and datasets. The PyTorch weight initialization is made for all the layers up to the last in the 7 models already implemented in PyTorch. The other 3 models are always randomly initialized with a fixed seed.
Hyperparameter Searched Values
Initial Learning Rate ($\text{lr}_{0}$) $1,2,3,4,5,6,7,8,9,10]\times 10^{-4}$
Weight Initialization ($\theta_{0}$) PyTorch, Random
Weight Decay ($\lambda$) Uniform(0, $10^{-3}$)
Batch Size ($B$) 64, 128
Optimizer (Op.) SGD, Adam [58], AdamW [59]
Dropout (p) 0.0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7
4.1.3 Evaluation Metrics
A large variety of metrics have been proposed by the solar forecasting community in order to assess the quality of a model’s predictions in the most insightful way possible [60, 61, 62]. In this study, we evaluate the root mean squared error (RMSE), the mean absolute error (MAE) and the normalized RMSE (nRMSE), which are well established metrics for evaluating the overall fit of the predictions:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(I_{i}-\hat{I}_{i})^{2}}$ (14)
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|I_{i}-\hat{I}_{i}|$ (15)
$\text{nRMSE}=\frac{\text{RMSE}}{\frac{1}{N}\sum_{i=1}^{N}I_{i}}$ (16)
where $N$ is the number of samples, $I_{i}$ is the ground truth measurement of the GHI for the $i^{th}$ sample and $\hat{I}_{i}$ is the model’s estimate of the GHI for the $i^{th}$ sample. We employ the RMSE as our main ranking metric, since it penalizes larger errors more heavily than the MAE. The nRMSE is used specifically to analyze performance across different weather seasons and hours of the day, due to the scale-invariant nature of this metric.
Note that, regardless of our choice for the model’s target variable, the evaluation metrics are always expressed in terms of the GHI ($\text{W/m}^{2}$) or as a percentage of GHI (in the case of nRMSE), as that is the variable of interest.
4.2 Forecasting
In the forecasting task, we use $\mathcal{D}^{\text{sirta}}$ with the goal of generating future GHI values with lead times ranging from 2 to 10 minutes, based on 8 minutes of past ASI context. We investigate two approaches: a single-step approach, where an irradiance forecasting module directly predicts the future irradiance values, and a two-step approach, where a video prediction module first generates future ASI frames, and a subsequent module estimates the corresponding irradiance values for each generated frame. This irradiance estimation module corresponds to our best GHI estimation module, which is trained on the ground truth images.
4.2.1 Data Processing
In order to create the ASI sequence samples, we closely followed the data processing pipeline in [40], which is used as the baseline for our forecasting results. This lead to a few modifications to the data processing steps used in our GHI estimation experiments, described in section 3. Briefly, the key differences are: 1) the test set only includes samples that correspond to the odd days of 2019; 2) each image in the input sequence has a $128\times 128$ resolution.
Regarding the use of only odd days of 2019 as the test set, this only means that the test set in the GHI forecasting experiments will be about half the size of the test set in the GHI estimation experiments. We note, however, that the training set years remain the same, which means that, for the two-step approach, there is no data leakage between the video prediction and the GHI estimation modules.
Regarding the use of $128\times 128$ image resolution, this means that we need to resize the generated images to $64\times 64$ prior to the GHI estimation module in the two-step approach.
4.2.2 Two-step Approach
Given an input batch of image sequences $X_{i}\in\mathbb{R}^{B\times T\times C\times H\times W}$, the model is trained to predict the output image sequence $Y_{i}\in\mathbb{R}^{B\times T\times C\times H\times W}$ by minimizing the pixel-wise MSE, where $B$ is the batch size, $T$ is the number of frames, $C$ is the number of image channels, $H$ is the image height and $W$ is the image width. We use $T=5$ frames, such that the input consists of frames at times $\{t-8,t-6,t-4,t-2,t\}$, and the model predicts frames at $\{t+2,t+4,t+6,t+8,t+10\}$, in minutes. Our batch size is set to 16.
We employ an encoder-only Vision Transformer, closely following the design of PredFormer [63], which transforms the input image sequence $X_{i}$ into a sequence of $T\cdot N$ non-overlapping patches, where $N=\frac{HW}{p^{2}}$ is the number of patches per frame and $p$ is the patch size. We use a patch size of 16, resulting in $N=64$ patches per frame. Each image patch gets transformed into a patch embedding of dimension $D=256$, resulting in a $X_{p}\in\mathbb{R}^{B\times T\times N\times D}$ tensor.
A positional encoding is then added to $X_{p}$ to provide both the spatial and temporal information of each patch. Following this, each patch embedding will go through the stack of PredFormer layers, before finally being decoded back into the image patches by the linear patch decoder layer, which decodes $X_{p}$ into $Y_{i}$.
Regarding the PredFormer layers, we use full spatial-temporal self-attention, which allows each query patch to attend to all other patch tokens across the input sequence. Under this setup, each query patch will predict the patch at the exact same spatial position, but 10 minutes into the future, as illustrated by Figure 11. For more details on the model architecture, we refer the reader to the PredFormer article [63].
Finally, to obtain the GHI predictions, each of the $T$ predicted frames in $\hat{Y}_{i}$ are resized to a $64\times 64$ resolution, and are then fed to our best pre-trained estimation model to obtain the corresponding GHI estimate.
Figure 11: Input patches (left) and output patches (right) of the vision transformer model. Each query patch (red) can attend to all input patch tokens through full spatial temporal self-attention (blue). Each patch has size $16\times 16$, however, for better visualization, we show patches of size $32\times 32$ above.
4.2.3 Single-step Approach
This approach is conceptually similar to the first step of the two-step approach, and also follows the design of Predformer. However, instead of training the model to predict the future image frames, it is trained to directly predict the output clear sky index sequence $k_{t,i}\in\mathbb{R}^{B\times T}$ by minimizing the MSE
$L_{\text{MSE}}=\frac{1}{B\cdot T}\sum_{i=1}^{B}\sum_{t=1}^{T}\left(k_{t,i}-\hat{k}_{t,i}\right)^{2}$ (17)
where $B$ and $T$ have the exact same values used in the first step of the two-step approach. To optimize the PredFormer model with this loss function, two other changes were made. First, we prepend a learnable token embedding $e\in\mathbb{R}^{1\times 1\times 1\times D}$ to every frame of every batch sample in $X_{p}$, such that the new $X_{p}\in\mathbb{R}^{B\times T\times(N+1)\times D}$. This is analogous to the CLS token used in BERT [64]. Second, the linear patch decoder layer is replaced with a linear irradiance regression head. Again, akin to BERT, this linear regression layer is only applied to the output of the $e$ token, obtaining the clear sky index prediction, which we transform into a GHI estimate through Equation 2. Everything else is the same as in the first step of the two-step approach.
4.2.4 Evaluation Metrics
We evaluate the results at every timestep $\{t+2,t+4,t+6,t+8,t+10\}$ using both the RMSE and the forecasting skill (FS). The forecasting skill is a metric that measures the overall improvement over the SPM, according to:
$\text{FS}=\left(1-\frac{\text{RMSE}}{\text{RMSE}_{\text{SPM}}}\right)$ (18)
where $\text{RMSE}_{\text{SMP}}$ is the error of the SPM predictions, given by Equation 3. A positive FS implies that a model is better than the SPM and a negative value suggests otherwise.
5 Experiments and Results
5.1 Benchmarking Results
Table 5 shows the RMSE and MAE for the test set for all ten models and all three datasets, as well as the aggregated performance across all datasets. From these results, we observe that model depth and the usage of residual connections improves performance. For example, the ResNet50 and EfficientNet models, which are both deep residual models, consistently outperformed the others, especially under cloudy sky conditions. This is expected, since deeper networks are better equipped to extract complex spatial features, which are crucial under these sky conditions, due to the high variety of shapes, sizes and possible overirradiance effects that clouds can have.
However, models that are deep but lack residual connections, such as SolarNet and VGG16, showed significantly worse performance. This aligns with the findings of [10], who demonstrated that residual connections facilitate the information and gradient flow through the network, allowing for a much more efficient training of very deep models. The shallowest models in our benchmark, Sunset and UNet, yielded the worst performance. Their limited depth is likely a determinant factor in this poor performance, since it restricts their ability to extract the complex spatial features of clouds.
Finally, the results show that $\mathcal{D}^{\text{folsom}}$ achieved the best “All” performance, despite yielding worst results on both the “Clear” and “Cloudy” subsets (an instance of Simpson’s paradox). The best “All” performance can be attributed to the fact that over 53% of the samples in $\mathcal{D}^{\text{folsom}}$ are clear sky samples (as depicted in Table 3), which will naturally bias the overall error toward lower values. On the other hand, the performance on the other two subsets was much worse, even compared to $\mathcal{D}^{\text{nrel}}$, which has far fewer training samples. We hypothesize that this behavior is due to the lens contamination we highlighted in subsection 3.1 and Figure 3, since the model can be mistaking the dirt stains for false clouds, hurting the performance on both the sky conditions.
Table 5: Model architecture benchmark results for all three datasets. The best performing architecture is highlighted in green. The Aggr. dataset refers to the aggregated predictions/ground truth for all three datasets.
RMSE [W/m²] MAE [W/m²]
Sky cond. All Cloudy Clear All Cloudy Clear
Dataset
Folsom EfficientNet 37.39 ± 0.12 52.60 ± 0.18 13.86 ± 0.51 20.87 ± 0.25 32.75 ± 0.20 10.40 ± 0.32
MobileNetV2 45.17 ± 2.04 62.22 ± 2.11 20.60 ± 3.06 27.94 ± 1.93 41.67 ± 1.87 15.85 ± 2.20
RegNet 37.68 ± 0.36 53.21 ± 0.37 13.31 ± 0.79 20.91 ± 0.47 33.21 ± 0.39 10.09 ± 0.58
ResNet18 40.26 ± 0.81 56.31 ± 1.17 16.02 ± 0.57 22.92 ± 0.50 35.18 ± 0.84 12.13 ± 0.37
ResNet34 45.91 ± 11.74 54.34 ± 0.52 31.65 ± 24.91 22.84 ± 1.84 33.37 ± 0.17 13.57 ± 3.41
ResNet50 37.21 ± 0.26 52.62 ± 0.35 12.89 ± 0.39 20.58 ± 0.15 32.82 ± 0.25 9.80 ± 0.23
SolarNet 43.25 ± 1.81 59.87 ± 1.70 18.90 ± 2.93 25.32 ± 1.43 38.12 ± 1.13 14.06 ± 1.69
Sunset 48.08 ± 2.31 65.34 ± 2.64 24.06 ± 3.70 29.16 ± 2.18 41.89 ± 1.92 17.96 ± 2.85
UNet 50.12 ± 2.27 64.29 ± 1.45 32.67 ± 5.14 33.57 ± 2.89 43.10 ± 1.47 25.18 ± 4.38
VGG16 39.64 ± 0.52 55.35 ± 0.57 16.03 ± 0.78 22.91 ± 0.47 35.28 ± 0.46 12.02 ± 0.54
SIRTA EfficientNet 40.19 ± 0.23 44.00 ± 0.27 11.42 ± 0.42 23.72 ± 0.21 27.00 ± 0.31 8.53 ± 0.43
MobileNetV2 40.61 ± 0.27 44.44 ± 0.32 11.91 ± 1.00 23.56 ± 0.21 26.72 ± 0.22 8.94 ± 0.67
RegNet 45.25 ± 2.09 49.03 ± 1.88 19.29 ± 5.48 27.99 ± 1.70 31.05 ± 1.27 13.78 ± 3.74
ResNet18 44.80 ± 2.24 48.66 ± 2.41 18.21 ± 2.08 27.90 ± 1.74 31.14 ± 1.94 12.89 ± 1.22
ResNet34 43.72 ± 2.00 46.60 ± 1.22 25.32 ± 9.00 26.71 ± 1.47 28.95 ± 0.91 16.31 ± 4.28
ResNet50 39.68 ± 0.21 43.46 ± 0.22 10.75 ± 0.56 23.06 ± 0.16 26.30 ± 0.14 8.07 ± 0.42
SolarNet 42.99 ± 0.58 47.04 ± 0.63 12.57 ± 0.44 25.21 ± 0.43 28.64 ± 0.47 9.31 ± 0.37
Sunset 46.90 ± 1.03 50.97 ± 0.94 18.77 ± 2.81 29.46 ± 0.96 32.81 ± 0.80 13.94 ± 1.96
UNet 47.75 ± 1.44 51.31 ± 1.50 25.36 ± 1.72 30.68 ± 0.96 33.33 ± 0.92 18.43 ± 1.29
VGG16 52.57 ± 2.81 56.16 ± 2.50 30.62 ± 5.96 33.67 ± 2.30 36.30 ± 1.74 21.48 ± 4.92
NREL EfficientNet 42.44 ± 4.58 52.14 ± 1.64 19.21 ± 13.28 23.97 ± 1.25 33.43 ± 1.11 10.53 ± 1.57
MobileNetV2 47.18 ± 1.18 60.57 ± 1.49 13.31 ± 0.53 27.15 ± 0.58 39.36 ± 0.77 9.80 ± 0.33
RegNet 41.65 ± 0.33 53.01 ± 0.44 14.44 ± 0.49 24.42 ± 0.17 33.47 ± 0.19 11.58 ± 0.29
ResNet18 40.84 ± 0.57 52.36 ± 0.75 11.91 ± 1.76 23.17 ± 0.54 33.12 ± 0.37 9.03 ± 1.25
ResNet34 43.12 ± 3.10 53.56 ± 3.19 20.25 ± 4.87 26.05 ± 2.57 34.00 ± 2.42 14.77 ± 3.11
ResNet50 38.75 ± 0.29 49.85 ± 0.37 10.25 ± 0.58 21.73 ± 0.23 31.56 ± 0.38 7.77 ± 0.41
SolarNet 44.33 ± 0.83 56.84 ± 0.98 12.93 ± 1.07 25.58 ± 0.84 36.62 ± 0.79 9.92 ± 1.08
Sunset 55.47 ± 2.76 71.24 ± 3.46 15.46 ± 1.42 31.14 ± 1.30 45.40 ± 1.79 10.89 ± 0.66
UNet 52.21 ± 1.77 66.45 ± 2.30 18.05 ± 1.07 31.43 ± 0.96 44.35 ± 1.60 13.09 ± 0.87
VGG16 41.45 ± 0.95 53.19 ± 1.19 11.80 ± 0.81 23.62 ± 0.69 33.76 ± 0.73 9.24 ± 0.75
Aggregated EfficientNet 38.53 ± 0.31 49.07 ± 0.12 14.13 ± 1.48 21.89 ± 0.22 30.31 ± 0.19 10.17 ± 0.30
MobileNetV2 43.95 ± 1.34 55.19 ± 1.24 19.39 ± 2.71 26.57 ± 1.25 35.13 ± 0.99 14.66 ± 1.86
RegNet 40.33 ± 0.86 51.45 ± 0.90 14.37 ± 1.39 23.24 ± 0.73 32.29 ± 0.70 10.64 ± 0.83
ResNet18 41.74 ± 0.67 52.98 ± 0.88 16.17 ± 0.26 24.45 ± 0.46 33.34 ± 0.68 12.08 ± 0.22
ResNet34 45.41 ± 8.03 51.12 ± 0.46 31.40 ± 21.50 24.17 ± 1.30 31.50 ± 0.29 13.98 ± 3.00
ResNet50 38.05 ± 0.20 48.75 ± 0.24 12.51 ± 0.36 21.39 ± 0.14 29.95 ± 0.17 9.48 ± 0.22
SolarNet 43.23 ± 1.36 54.56 ± 1.21 17.96 ± 2.59 25.30 ± 1.06 33.97 ± 0.79 13.25 ± 1.45
Sunset 48.14 ± 1.35 59.94 ± 1.44 23.15 ± 3.06 29.35 ± 1.23 38.16 ± 0.86 17.10 ± 2.27
UNet 49.54 ± 1.35 59.18 ± 0.97 31.29 ± 4.41 32.58 ± 1.78 38.96 ± 0.79 23.71 ± 3.60
VGG16 44.08 ± 1.06 55.61 ± 1.11 18.50 ± 1.42 26.22 ± 0.78 35.64 ± 0.79 13.11 ± 0.79
Figure 12 summarizes the aggregated results in Table 5 for both the clear and cloudy sky test subsets.
Figure 12: Cloudy and clear sky scatter plot showing the aggregated performance of all ten models. The left panel shows the RMSE performance and the right panel shows the MAE. Marker sizes are scaled according to the number of FLOPs for each model.
5.2 Ablation Experiments
Based on the results from the previous section, we then perform several ablation experiments using the ResNet50 model with the goal of investigating the impact that different configurations of pre-processing steps and training setups have on the performance. These configurations are summarized in Table 6. Note that, for the $\Delta\text{t}$” and “Sample Interval” configuration, the searched values are only applied to the training set, while the default values are used for the test set.
Table 6: Configuration table for all the ablation experiments in this study. The values for a certain configuration are only searched in their respective sections, otherwise the default value is used.
Configuration Dataset Searched Values Default Values
Folsom File Name, Date Modified Date Modified
Image Timestamps (5.2.1) SIRTA - File Name
NREL - File Name
Folsom -30s:10s:30s -20s
Time off-set $\Delta\text{t}$ (5.2.2) SIRTA -30s:10s:30s 10s
NREL -30s:10s:30s -30s
Folsom
Model Target (5.2.3) SIRTA GHI, $k_{t}$, $K_{t}$, $k_{t}^{\text{weighted}}$, $K_{t}^{\text{weighted}}$ $k_{t}$
NREL
Folsom Long
Camera Exposure (5.2.4) SIRTA Short, Long Long
NREL Short, Long
Folsom 1 - 10 min 1 min
Sample Interval (5.2.5) SIRTA 1/2 - 10 min 1/2 min
NREL 10 min 10 min
Folsom
Sun Mask (5.2.6) SIRTA Mask, No Mask No Mask
NREL
5.2.1 Image Timestamps in $\mathcal{D}^{\text{folsom}}$
Table 7: GHI and DNI RMSE results for the four possible train/test combinations (FN = File Name; DM = Date Modified) and the different test subsets in $\mathcal{D}^{\text{folsom}}$.
RMSE $\text{W/m}^{2}$]
Sky Cond. Train Test GHI DNI
All FN FN 62.52 ± 0.18 96.04 ± 0.75
FN DM 56.72 ± 0.13 81.31 ± 1.28
DM FN 59.37 ± 0.34 96.60 ± 0.45
DM DM 37.21 ± 0.26 55.42 ± 0.94
Cloudy FN FN 89.76 ± 0.35 135.81 ± 0.94
FN DM 81.19 ± 0.18 113.88 ± 1.66
DM FN 85.34 ± 0.48 135.39 ± 0.36
DM DM 52.62 ± 0.35 73.25 ± 0.86
Clear FN FN 16.05 ± 0.60 33.13 ± 3.50
FN DM 15.76 ± 0.57 31.77 ± 3.49
DM FN 14.69 ± 0.43 37.55 ± 1.60
DM DM 12.89 ± 0.39 32.42 ± 1.67
As mentioned in subsection 3.6, aligning the sky images with the irradiance data in $\mathcal{D}^{\text{folsom}}$ using the image file names led to large label errors, while the date modified seemed much more aligned with the irradiance data. To the best of our knowledge, this has not yet been reported in the literature so, in order to ascertain that the date modified is truly the correct timestamp for labeling the images in $\mathcal{D}^{\text{folsom}}$, a more quantitative analysis is also needed. To this end, we train two different models, one trained on a file name aligned training set and another on a date modified aligned training set. Both of these models are then evaluated on file named aligned testing set and a date modified aligned testing set, totaling four different train/test image timestamp alignment combinations. All other configurations are kept at their default values, depicted in Table 6, and, for this particular experiment, we also evaluate the results on the DNI component, both to rule out the overirradiance hypothesis raised in subsection 3.6, as well as to offer a benchmark for future studies that use $\mathcal{D}^{\text{folsom}}$ to focus on this specific component.
Table 7 depicts the GHI and DNI results. One can see that, when the training set and the testing set are aligned with the date modified timestamp, the RMSE drops significantly, confirming the assumptions made in subsection 3.6. Perhaps unsurprisingly, this drop was much larger for the cloudy subset than the clear subset, since, for the majority of clear sky samples, a few minutes difference in alignment results in small GHI/DNI differences. On the other hand, under cloudy sky conditions, the sky image and irradiance alignment is much more critical.
Moreover, Table 7 also highlights the fact that evaluating with the date modified alignment is always better than with the file name alignment, even when training with the file name. The reason for this can be attributed to Figure 7, which shows that the file name alignment is much worse in 2016 than in 2014/2015. Since we selected 2014/2015 for training and 2016 for testing, this means that the model can still learn somewhat useful features when trained with the file name alignment.
5.2.2 Time Shifting Operation
As described in subsection 3.5, we shift the irradiance measurements of the training set by applying a timestamp offset of $\Delta\text{t}$. For each dataset, we optimize $\Delta\text{t}\in\{-30\text{s}:10\text{s}:30\text{s}\}$ by applying a 5-fold CV (as described in subsubsection 4.1.1) and minimizing the validation loss. We then apply the optimal $\Delta\text{t}$ to the full training set to train the model and evaluate the results on the testing set.
The CV results are depicted in Figure 13, which shows the mean validation set RMSE across all 5 folds, for each dataset. Interestingly, in $\mathcal{D}^{\text{folsom}}$ and $\mathcal{D}^{\text{nrel}}$, the validation RMSE decreases slightly for a negative $\Delta\text{t}$ and increases for a positive $\Delta\text{t}$, but in $\mathcal{D}^{\text{sirta}}$ the opposite happens. This suggests that the type of average applied to the GHI samples in $\mathcal{D}^{\text{folsom}}$ was the same as in $\mathcal{D}^{\text{nrel}}$, but different in $\mathcal{D}^{\text{sirta}}$. The optimal $\Delta\text{t}$ was -20s, 10s and -30s for $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively.
Regarding the results of the test set, Table 8 shows that using the optimal $\Delta\text{t}$ to train the models always leads to a consistent gain in performance for all datasets, though in $\mathcal{D}^{\text{nrel}}$ this gain was close to negligible. The improvement is more apparent under cloudy sky conditions, since the sky image and irradiance alignment is more critical during these periods.
Figure 13: Validation set results for $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$. For each $\Delta\text{t}$, we report the mean RMSE across all 5 folds, as well as the standard deviation.
Table 8: Test set performance comparison between a 0 second $\Delta\text{t}$ (baseline) and the best performing one, obtained during CV.
RMSE [W/m²]
Dataset $\Delta\text{t}$ All Clear Cloudy
Folsom 0s 40.24 ± 0.94 15.60 ± 1.36 56.41 ± 1.06
-20s 37.21 ± 0.26 12.89 ± 0.39 52.62 ± 0.35
SIRTA 0s 41.22 ± 0.54 11.17 ± 0.44 45.15 ± 0.59
10s 39.68 ± 0.21 10.75 ± 0.56 43.46 ± 0.22
NREL 0s 39.01 ± 0.37 10.28 ± 0.53 50.20 ± 0.50
-30s 38.75 ± 0.29 10.25 ± 0.58 49.85 ± 0.37
5.2.3 Target Variable and Loss Function
Table 9: Test set results for all 5 different model target variables. The best performing target is highlighted in green for each dataset.
RMSE [W/m²] MAE [W/m²]
Sky cond. All Cloudy Clear All Cloudy Clear
Dataset
Folsom GHI 65.77 ± 2.22 82.62 ± 2.81 46.09 ± 1.80 48.24 ± 2.22 57.44 ± 2.29 40.14 ± 2.38
$k_{t}$ 37.21 ± 0.26 52.62 ± 0.35 12.89 ± 0.39 20.58 ± 0.15 32.82 ± 0.25 9.80 ± 0.23
$K_{t}$ 38.16 ± 0.62 53.09 ± 0.76 16.02 ± 0.90 22.09 ± 0.60 33.42 ± 0.70 12.12 ± 0.63
$k_{t}$ (weighted) 37.46 ± 0.15 52.87 ± 0.21 13.35 ± 0.45 21.02 ± 0.20 33.33 ± 0.32 10.19 ± 0.36
$K_{t}$ (weighted) 38.93 ± 0.66 53.87 ± 0.74 17.17 ± 0.89 23.06 ± 0.69 34.39 ± 0.71 13.08 ± 0.75
SIRTA GHI 50.84 ± 2.18 54.29 ± 2.24 29.95 ± 3.10 34.69 ± 1.92 36.93 ± 1.91 24.30 ± 2.80
$k_{t}$ 39.68 ± 0.21 43.46 ± 0.22 10.75 ± 0.56 23.06 ± 0.16 26.30 ± 0.14 8.07 ± 0.42
$K_{t}$ 39.33 ± 0.43 42.91 ± 0.42 13.39 ± 1.14 23.13 ± 0.45 25.89 ± 0.35 10.33 ± 0.99
$k_{t}$ (weighted) 39.88 ± 0.31 43.72 ± 0.35 10.22 ± 0.42 23.34 ± 0.24 26.70 ± 0.31 7.78 ± 0.27
$K_{t}$ (weighted) 39.23 ± 0.28 42.94 ± 0.30 11.30 ± 0.80 22.97 ± 0.28 26.04 ± 0.26 8.71 ± 0.73
NREL GHI 42.43 ± 0.68 54.12 ± 0.90 14.04 ± 0.37 25.94 ± 0.36 36.64 ± 0.58 10.74 ± 0.20
$k_{t}$ 38.75 ± 0.29 49.85 ± 0.37 10.25 ± 0.58 21.73 ± 0.23 31.56 ± 0.38 7.77 ± 0.41
$K_{t}$ 38.20 ± 0.24 48.91 ± 0.36 11.60 ± 0.48 21.97 ± 0.11 31.16 ± 0.19 8.91 ± 0.40
$k_{t}$ (weighted) 38.06 ± 0.33 48.94 ± 0.43 10.18 ± 0.51 21.45 ± 0.30 31.11 ± 0.41 7.72 ± 0.25
$K_{t}$ (weighted) 38.04 ± 0.26 48.78 ± 0.35 11.13 ± 0.18 21.83 ± 0.28 31.15 ± 0.36 8.60 ± 0.19
In this experiment, we consider five combinations of target variables and loss functions. Specifically, for the MSE loss (12), we consider three choices for the target variable $y\in\{I,k_{t},K_{t}\}$, resulting in the following expressions for the GHI estimate:
$\hat{I}_{i}=\begin{cases}\hat{y}_{i},&\text{if $y_{i}=I_{i}$},\\ I^{\text{clr}}\hat{k}_{t,i},\;\text{where $\hat{k}_{t,i}=\hat{y}_{i}$},&\text{if $y_{i}=k_{t,i}$},\\ I^{\text{extr}}\hat{K}_{t,i},\;\text{where $\hat{K}_{t,i}=\hat{y}_{i}$},&\text{if $y_{i}=K_{t,i}$}.\end{cases}$ (19)
[…]
While the use of $k_{t}$ or $K_{t}$ as target solves the seasonal variance problem, it fails to take into account the diurnal variance of the GHI, i.e. a sample at the end of the day (with low GHI magnitude) will be weighted the same as a sample in the middle of the day (with high GHI magnitude). This could hurt the overall performance of the model, as it reduces the emphasis on middle-of-the-day samples, which are typically associated with larger prediction errors (RMSE and MAE). To solve this, we rescale the model prediction back into the GHI estimate before computing the squared error, leading to the loss function:
$L_{\text{MSE}}^{\text{weighted}}=\frac{1}{B}\sum_{i=1}^{B}\left(I^{\text{clr/extr}}(y_{i}-\hat{y}_{i})\right)^{2}=\frac{1}{B}\sum_{i=1}^{B}(I_{i}-\hat{I}_{i})^{2}$ (20)
where
$I^{\text{clr/extr}}=\begin{cases}I^{\text{clr}},&\text{if $y_{i}=k_{t,i}$},\\ I^{\text{extr}},&\text{if $y_{i}=K_{t,i}$}\end{cases}$ (21)
and $\hat{I}_{i}$ is given by (19). Note that this is equivalent to the solution presented in [42].
Figure 14: GHI and $k_{t}$ values comparison for a clear sky day showing the seasonal variance in $\mathcal{D}^{\text{folsom}}$. The upper panel shows the GHI values and the lower panel shows the $k_{t}$ values. The corresponding images to each labeled marker are displayed on the right.
The results for this experiment are summarized in Table 9, from which it is clear that modeling the raw GHI values yields the worst results by a large margin. This is especially true for $\mathcal{D}^{\text{folsom}}$, where there is almost a 50% improvement when using other target variables. However, regarding the other target variables, the performance was largely the same, even when using the weighted loss function, which suggests that weighting midday samples more heavily may not be as beneficial as was initially hypothesized. Thus, we opt to use $k_{t}$ as our default target variable.
We also compare model performance across seasons and hours of the day, using either GHI or $k_{t}$ as target variable, and depict the results in Figure 15 and Figure 16. Since both the season and hour of day affect the magnitude of the GHI values, scale-dependent metric comparisons (such as RMSE and MAE) become more difficult. Consequently, we make use of the nRMSE metric for this analysis, due to its scale-invariant nature.
From Figure 15 one can see that using $k_{t}$ over GHI as the target variable yields a consistent gain in performance across all seasons in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, while in $\mathcal{D}^{\text{folsom}}$ this gain was slightly larger in spring and winter. We also note that the model performance tended to be best during summer and worse in winter. Since the summer season typically contains the highest number of clear sky samples, GHI estimation becomes much easier during this season. Additionally, extended daylight hours during summer increase the number of available sky images, which may introduce a slight model bias towards this season.
Regarding the hourly results depicted in Figure 16, we observe that models trained with $k_{t}$ target consistently outperform those trained with GHI throughout the day, with particularly notable improvements around sunrise and sunset. This is significant, as these periods typically coincide with large net load ramps, where accurate irradiance predictions are critical for efficient grid management.
Figure 15: Seasonal nRMSE for the GHI and $k_{t}$ target variables.
Figure 16: Hourly nRMSE for the GHI and $k_{t}$ target variables. For each hour mark, the nRMSE is computed across all samples between zero and 59 minutes.
5.2.4 Camera Exposure
As shown in Figure 2, the images in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ provide two different camera exposure levels, a long exposure form and a short exposure form. Although we use the long exposure form as our default value for all other experiments, the short exposure form provides more detail in the Sun disk area at the cost of providing less detail everywhere else. We hypothesize that in the GHI estimation task, predicting the irradiance value given a short exposure image is a much easier task, since the main region of interest is the area around the Sun. To test this hypothesis, we train and evaluate models using the short exposure images in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ and compare the results to their long exposure counterparts.
From the results presented in Table 10, using the short exposure images improved the performance on both datasets, especially under cloudy sky conditions. This behavior is expected, since the long exposure images can have saturated pixels around that Sun that make it hard to discern between a Sun and a cloud pixel. Although this confirms our hypothesis that the short exposure images yield better results on the GHI estimation task, we have not tested it on the forecasting task, which is why we default to the long exposure images in our other experiments.
Table 10: Test set results using the two camera exposures available in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$.
RMSE [W/m²] MAE [W/m²]
Exposure All Cloudy Clear All Cloudy Clear
Dataset
SIRTA Long 39.68 ± 0.21 43.46 ± 0.22 10.75 ± 0.56 23.06 ± 0.16 26.30 ± 0.14 8.07 ± 0.42
Short 36.46 ± 0.31 39.86 ± 0.34 11.19 ± 0.45 21.62 ± 0.15 24.58 ± 0.15 7.93 ± 0.31
NREL Long 38.75 ± 0.29 49.85 ± 0.37 10.25 ± 0.58 21.73 ± 0.23 31.56 ± 0.38 7.77 ± 0.41
Short 37.34 ± 0.45 48.03 ± 0.61 9.86 ± 0.30 21.01 ± 0.22 30.52 ± 0.35 7.52 ± 0.32
Table 11: Sun mask results for the testing set in all three datasets. Note that the approach of [22] uses 2014 in $\mathcal{D}^{\text{folsom}}$ as the testing set, but in our approach we use 2016.
RMSE [W/m²] MAE [W/m²]
Target GHI $k_{t}$ GHI $k_{t}$
Dataset
Folsom No Mask [22] 64.83 - 37.23 -
Mask [22] 60.31 - 36.02 -
No Mask [Ours] 65.77 ± 2.22 37.21 ± 0.26 48.24 ± 2.22 20.58 ± 0.15
Mask [Ours] 56.60 ± 4.95 39.90 ± 0.52 41.18 ± 5.13 22.79 ± 0.40
SIRTA No Mask 50.84 ± 2.18 39.68 ± 0.21 34.69 ± 1.92 23.06 ± 0.16
Mask 43.67 ± 1.51 40.83 ± 0.74 27.31 ± 1.45 24.04 ± 0.69
NREL No Mask 42.43 ± 0.68 38.75 ± 0.29 25.94 ± 0.36 21.73 ± 0.23
Mask 42.98 ± 0.52 41.62 ± 0.73 25.48 ± 0.40 22.83 ± 0.34
5.2.5 Sample Interval
In this experiment we change the size of the training set by increasing the sample interval from 1 to 10 minutes. Since in $\mathcal{D}^{\text{nrel}}$ the minimum sample interval is 10 minutes, we do not search for any sample interval values using this dataset. All other configurations are kept at their default values.
The results depicted in Figure 17 show that the test error rapidly increases with larger sample intervals under cloudy sky conditions, whereas under clear sky conditions the increase is much smaller. This behavior is expected since modeling clear sky conditions is a relatively simple task, simple enough that little new information would be gained by providing the model with additional clear sky training examples. In contrast, cloudy sky conditions can exhibit high variability due to complex cloud dynamics, shapes and overirradiance effects. As a result, decreasing the number of training examples rapidly decreases the model’s robustness under these conditions.
Additionally, it is also interesting to note that, when all datasets are on “equal ground” with a 10 minute sample interval, the behavior mentioned in subsection 5.1 is even more noticeable, that is, the “Cloudy” test set performance in $\mathcal{D}^{\text{folsom}}$ gets significantly worse by comparison. This further validates our hypothesis that the model is mistaking the lens contamination stains for false clouds.
Figure 17: Test set RMSE for different sample intervals applied only to the training set. Only a single marker is shown for $\mathcal{D}^{\text{nrel}}$ since the minimum sample interval in this dataset is 10 minutes. The 1 minute sample interval value for $\mathcal{D}^{\text{sirta}}$ is only for the 2017 training year, since in 2018 the minimum sample value is 2 minutes.
5.2.6 Sun Mask
In the Sun mask experiment, we compare the model performance for 4 different combinations: 1) GHI target with no mask; 2) $k_{t}$ target with no mask; 3) GHI target with mask; 4) $k_{t}$ target with mask. In 3) and 4), we append the Sun mask described in subsection 3.7 as a fourth channel to the RGB input image. This implies that the first convolutional layer of the model needs to be redesigned such that it accepts a 4-channel input, which also means that the kernel weights corresponding to this additional channel need to randomly initialized, since no pre-trained weights are available. All other configurations are kept at their default values presented in Table 6.
In Table 11 we show the GHI estimation results for the different mask/target combinations, and compare it to [22]. Even though they used a different testing year for $\mathcal{D}^{\text{folsom}}$, we were also able to confirm, for $\mathcal{D}^{\text{folsom}}$ and $\mathcal{D}^{\text{sirta}}$, that appending a Sun mask as a fourth channel to the input improves the model performance when the GHI is its target, while doing so in $\mathcal{D}^{\text{nrel}}$ seemed to yield negligible results. However, when the $k_{t}$ is the target, the Sun mask actually degrades the model performance in all datasets. We hypothesize that this is because of the fact that the Sun position in the image is less relevant when modeling the $k_{t}$, as discussed in subsubsection 5.2.3, such that appending a Sun mask would mostly add redundant information to the model, which could increase the risk of over-fitting the training set. Another possible explanation is the necessary redesign of the model to accept a 4-channel input, which forced a random weight initialization in the added weights of the first layer of the model. Since all other weights are initialized with their pre-trained values, this can disrupt the pre-learned low-level features extracted by the first layer.
Figure 18: Ground truth and GHI predictions from the two-step and single-step approaches on two distinct future time windows: one where the Sun transitions from covered to visible (top panel), and the other where it transitions from visible to covered (bottom panel). In each panel, the left section displays the ASI context used as input for the forecasting model. The right section displays: ground truth future ASI (top row); ground truth and predicted future GHI values (middle row); predicted future ASI used as input for the GHI estimation model in the two-step approach (bottom row).
Table 12: Forecasting results for all models across all lead times. The proposed two-step approach showed significant improvements over the single-step approach, as well as the state of the art ECLIPSE [40] model. A lower bound on the forecasting error of the two-step approach is obtained by estimating the GHI using the ground truth instead of the predicted sky images.
RMSE [W/m2] (Forecast Skill [%])
2 min 4 min 6 min 8 min 10 min
Smart Persistence 93.6 (0.0%) 117.4 (0.0%) 129.8 (0.0%) 137.6 (0.0%) 143.1 (0.0%)
ECLIPSE [40] 83.8 (10.2%) - 98.5 (23.6%) - 109.1 (24.0%)
Single-step approach 90.5 (3.3%) 102.2 (13.0%) 108.2 (16.7%) 111.1 (19.3%) 113.9 (20.4%)
Two-step approach (Proposed) 80.9 (13.6%) 88.8 (24.3%) 94.6 (27.2%) 99.8 (27.5%) 103.9 (27.4%)
GHI estimation from real future ASI 44.1 (52.9%) 44.1 (62.4%) 44.1 (66.0%) 44.1 (67.9%) 44.1 (69.2%)
5.3 Forecasting Results
We begin our forecasting analysis by comparing the predictions from the two-step and single-step approaches on two distinct time windows. The first (top panel of Figure 18), where the Sun transitions from covered to visible, corresponds to the window with the highest standard deviation in ground truth future GHI values. The second (bottom panel of Figure 18), where the Sun transitions from visible to covered, corresponds to the window with the highest standard deviation in predicted future GHI values. Both cases show that the single-step approach tends to avoid large errors by simply predicting a mean GHI value, which comes at the cost of failing to predict GHI ramp events. In contrast, the two-step approach is capable of identifying such events, albeit with some limitations. By decoupling future event prediction from GHI estimation, the model is able to capture the cloud dynamics more effectively, enabling it to correctly predict when the Sun will be covered and when it will be visible. Consequently, when the generated images are combined with a GHI estimation model, the two-step approach yields much more accurate and sensible predictions.
However, we note that the generated images are far from realistic, exhibiting noticeable blur and heavy artifacts along the patch borders. The blur is likely due to our choice of loss function, which was only the pixel-wise MSE, while the border artifacts likely stems from our choice of linear decoder, which independently decodes each patch embedding. Nevertheless, we believe these results provide a strong proof of concept for the two-step approach in GHI nowcasting.
In Table 12 we compare the test set results for all forecasting models. These results further corroborate our proposed two-step approach, since it shows a significant improvement over the other models across all lead times. However, the lower bound forecasting error, which we obtain by replacing the predicted sky images with the ground truth before the GHI estimation module, indicates that there is still a large performance gap to be gained by improving the forecasting module and generating more realistic images.
6 Conclusion
Due to the intermittent nature of solar PV energy, the growing integration of this energy source into the electrical grid presents several technical challenges, such as optimizing the operational cost of the grid and maintaining its frequency and voltage stability. To address this issue, recent studies have applied deep learning-based architectures and methods to forecast the short-term irradiance (i.e. nowcasting) from all-sky images. In this study, we approach the nowcasting problem using a two-step framework: first we predict the future ASI with a video prediction module; second we estimate the corresponding GHI to each predicted future image frame with an irradiance estimation module. Our primary focus is on the GHI estimation module, and we begin our analysis with a benchmark of 10 different model architectures. This benchmark revealed that the ResNet50 model was the best performer across all datasets and highlighted that model depth and residual connections can improve model performance, especially under cloudy sky conditions.
Using the ResNet50 model, we then conduct ablation experiments and uncover several key findings: 1) in the Folsom dataset, the image file names are misaligned with the irradiance timestamps, but the correct alignment can be achieved by using the image date modified metadata; 2) shifting the irradiance measurements by a small $\Delta\text{t}$ during training improves performance by mitigating the backward/forward averaging effects applied to these measurements; 3) predicting the clear-sky index $k_{t}$ yields substantial performance gains compared to predicting GHI directly; 4) appending a Sun mask as a fourth image channel only benefits the model performance when predicting the GHI directly; among other insights.
We conclude our study by combining our best GHI estimation model with a video prediction model, achieving state of the art performance on the SIRTA dataset and providing strong proof of concept results for this two-step approach. These results highlight the potential for significant advancements in both the video prediction and GHI estimation modules, and we hope they inspire further research using the two-step approach in the GHI nowcasting literature.
Declaration of Competing Interest
The authors declare that they have no competing interests.
Data Availability
All three datasets used in this study are publicly available online. The Folsom dataset [32] is available at https://doi.org/10.5281/zenodo.2826939. The NREL dataset [46] is available at https://doi.org/10.7799/1052221. The SIRTA dataset [39] can be requested by filling out a data request form, which is to be found at the SIRTA observatory website.
Acknowledgements
The authors would like to acknowledge the SIRTA observatory for kindly providing the sky image and irradiance data used in this study. The authors would also like to acknowledge the authors of the Folsom and NREL datasets, for open-sourcing their respective datasets.
The work of D. Silva was supported in part by the Brazilian National Council for Scientific and Technological Development (CNPq-Brazil) under Grant 304619/2022-1.
The work of L. Varaschin was supported in part by Fundacao de Amparo à Pesquisa e Inovacao do Estado de Santa Catarina (FAPESC), Edital 61/2024.
This work was supported in part by computational resources funded by FAPESC under Grant 2024TR000090.